OpenStudy (anonymous):

James covers a distance of 8 meters in the first second of a 100 meter race. Then, he sprints at a speed of 9 meters per second. How far is James from the finish line 9 seconds after the race has started?

OpenStudy (hoblos):

in the first second he covered 8m then in the 8 more seconds he covered 8*9=72m so he totally covered 72+8=80m the distance left is 100-80=20m thus he is 20m far from the finish line

OpenStudy (anonymous):

Thanks : D

OpenStudy (hoblos):

my pleasure :D